Project

General

Profile

NDOS Commissioning Meeting 2011/07/11

Attendance: Jon, Susan, Raphael, Kirk, Alena, Leon, Sue, Kurt, Mat, Andrew, Carl, Martin

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • An additional ~30 golden APDs from dcm 3-1-2, 3-1-3, and 3-2-2 were added to the cooling list on Friday.
  • We lost an another 3 golden cooled APDs after the DSO scan this morning. Total lost is 11 out of 45 golden APDs.
  • We will enable golden APDs on the top of diblock 1 and 2 tomorrow morning.
  • NuMI target is down (water leak in the target, and the helium flow rate was maxed out). Next step is to install NT02 which should be able to withstand full power until the next target is available for installation (early September)

DAQ Status & Plans

  • Ron is out of town on vacation so development of not needing to reprogram the FEBs at the ConfigureHardware stage.
  • Sue: ability to disable the reprogram the FEBs is there, but the system has not shown to work if this doesn't happen.
  • Sue: new version of FEB firmware, version 18. First new version in 6 months; includes new feature that allows one to check that nanoslices are within microslice time window boundary. Time markers are produced at regular intervals, but backlog of data could cause nanoslices to fall outside time window.
  • Jon: what happens if/when a nanoslice is found to be outside window? Is the data dropped? Sue: we can test this on the teststand
  • Denis: is there a performance hit? Sue: not sure, again, we can find out from test.
  • Andrew: around 3 pm on Sunday, we lost connection to our DAQ cluster. Tracked down to an issue in GCC. The issue was resolved (not clear what was done), and Andrew brought the system back up. Required
  • Mat: was called at 1 am, novadaq-ctrl-runcontrol was down and he had to bring it back up.
  • Andrew: FNAL CD is not sure what happened
  • Andrew: we do not have 24/7 support in LCC. Jon: we should probably try to get that.
  • Discussion ensued about what qualifies as a "critical issue".

DCS Status & Plans

  • Martin: had meeting last Friday:
    • Gennadiy is looking into how to implement alarms
    • Jon: what's the difference between using IOC and "alarmer" package? Martin: "alarmer" is more capable of dealing with several small alarms, rather than writing complicated if-statements in the GUI.
    • Martin is working on overview "view" in GUI so that we don't have to have
  • Mat: what do we do if an APD goes into alarm? Jon:
  • Susan: noticed that ~3 FEBs are reporting ~-10C on APDs that are not being cooled. Martin will look into it.
  • Denis: do we know how to draw temp. vs. time? Martin: not yet, a work in progress.
  • Leon: there is a "pvmonitor" tool, will follow up with details and add it

AOB

  • Mat: can we work in the detector hall? Jon: by all means, let's make the best use of the beam downtime.